Agenda for 5/2/04 Board Meeting

Posted On April 29, 2004

OCA Board Meeting

May 2, 2004



  • Approval of March Minutes
  • Financial Summary/insurance status
  • Set July meeting date - BSC star party is 7/10, Anza 7/17.  7/4 and 7/25 look best on calendar.


  • Pad Coordinator issues: Stephen Eubanks asked that this be put on the May agenda at the end of the March meeting.  Bruce Crowe and Barbara Toy have been researching a number of legal issues relating to the pads that could affect our approach to the Pad Coordinator and any related positions, but have not been able to complete their analysis and recommendations.  They feel that it is premature for the Board to consider these issues at the May meeting and recommend that they be moved to the July agenda.  Approval as a Consent item would move consideration of this set of issues to the July meeting.


  • None


  • Repair/rebuilding of storage/library cabinets at Chapman urgently needed.
  • Anza site plan/Anza Planning Task Force
  • Anza Broadband - Implementation of payment schedule, controlling access, proposed rules for use of the service.
  • Observatory roof
  • Chapman University - show of appreciation.
  • RTMC
  • Kuhn - what can we do to move the repair/renovation project to completion?


  • Repairs needed to 12-inch LX200.
  • Star Member status for John Hoot
  • Anza Cleanup Day needed (clearing of rubbish as well as weed clearance)
  • Banquet
  • Posting pad licensee contact information at Anza site.
  • Moving of wires for weather station (which currently run across the computer room/control room at Anza House and pose a hazard to anyone walking across the room).


  • Club Archives:  We have an experienced archivist, Jon Bearscove, who has volunteered to become the club archivist and to set up the archive.
  • EOA/MOCAT:  Pier has been transported back to Anza, installed on the mounting bolts, and will be aligned when the 12" LX200 can be installed on it.  The computer from Buchheim has been set up in the MOCAT Observatory.  Tests that were run on the 12" LX200 to determine its tracking accuracy at the March star party found some significant problems that may be due to buildup on the gears; tentative arrangements were made to have the gears cleaned and relubedby Matt at Scope City, who volunteered to do it without charge.
  • Telescope Loaner Program - Bob Bell has taken a number of items out of the storage units to expand the equipment available in the Loaner Program, including the two Dobsonians that belonged to Don French and a small refractor.  Irvine Valley College also donated a tripod mount and a Vixen parallel mount that they were no longer using for the Loaner Program.
  • Items in storage - We still need to have the radio equipment evaluated to see if any has value and determine what should be done with it.  Barbara Toy took the boxes of smaller items to get them cleaned up and sorted for possible sale at RTMC.
  • Anza House - Larry Carr installed the vinyl flooring in the two west bathrooms at the April non-star party, along with the new toilets in those bathrooms.  The remaining bathroom should be done at the May star party.
  • Painting of Anza House
  • Mice/rat eradication in Anza House and observatory
  • Awards plans/program update
  • Tessmann update